A man is 25 years old when his daughter is born. How many years later will the daughter's age be half the father's age?
Answer by mananth(16946)

daughter = 0
father = 25
..
after x years
father = 25 +x
daughter = x
..
x=1/2(25+x)
multiply by 2
2x= 25+x
2x-x= 25
x= 25
..
after 25 years father will be double the daughter's age
..
m.annath@hotmail.ca